OpenStudy (anonymous):

factoring : 8x^6+7x^3-1

OpenStudy (anonymous):

(1+x) (2 x-1) (1-x+x^2) (1+2 x+4 x^2)

OpenStudy (anonymous):

can you teach me how did you get that answer?

OpenStudy (anonymous):

let t=x^3

OpenStudy (anonymous):

it is (t+1)(8t-1) then use A^3+B^3=(A+B)(A^2-AB+B^2)

OpenStudy (anonymous):

what do you call that method?

OpenStudy (anonymous):

A^3+B^3=(A+B)(A^2-AB+B^2) is a theorem..i dont know what the method is called

OpenStudy (anonymous):

ah it's called the difference of two cubes :D now i remember :D thanks :)